OpenStudy (anonymous):

Express answer in exact form. A regular hexagon with sides of 3" is inscribed in a circle. Find the area of a segment formed by a side of the hexagon and the circle. (Hint: remember Corollary 1--the area of an equilateral triangle is 1/4 s2 √3.)

OpenStudy (anonymous):

In the hint part, it's s^2. sorry about that.

jimthompson5910 (jim_thompson5910):

Area of one equilateral triangle: A = (1/4)*s^2*sqrt(3) A = (1/4)*3^2*sqrt(3) A = (1/4)*9*sqrt(3) A = (9/4)*sqrt(3) -------------------------------------------- The radius is equal to the side length of the equilateral triangle (drawing a pic really helps see this), so r = 3 So the area of the circle is A = pi*r^2 A = pi*(3)^2 A = 9pi So the area one-sixth of the circle is A = 9pi/6 = 3pi/2 Therefore, the answer is 3pi/2 - (9/4)*sqrt(3)